ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

10. Partial Differential Equations

1991 MS Classification: 35, 42, 46 Hasegawa, Kenji, Kogakuin University, Japan Boundedness of Pseudodifferential Operators Let a(x, D) be the pseudodifferential operator whose symbol is denoted by a(x, ξ). We shall present some results for boundedness of a(x, D) on various function spaces in the case that a(x, ξ) is Hölder continuous. Here we shall apply the theory of wavelets to decompose a(x, ξ).